Recently, More Money 4U offered an annuity that pays 5.7% compounded monthly. If $1,189 is deposited into this annuity every month, how much is in the account after 9 years? How much of this is interest?

Type the amount in the account: $____

Type the amount of interest earned: $____

Please do not round any numbers.

(Please square, label, or circle the answer)
Answer #1

Total amount after 9 years = $ 167,277.83

Total Interest Earned = $ 38,865.8

Working Notes:

Interest per month = 5.7%/12 = 0.475% = 0.00475

Future Value of Annuity = Annuity * ([1 + Interest] ^ N - 1 ) / Interest
= 1189 * ( [1+0.00475] ^108 - 1) / 0.00475)
= 167,277.83
Total Principal Paid = 1189 * 108
= 128,412.00
Interest Earned =167,277.83 - 128,412.0
= 38865.83